OpenStudy (anonymous):

how would you evaluate 3a^2-2ab+c if a=-1 b=0 and c=11 please show work.

OpenStudy (mimi_x3):

Just substitute a, b, and c in. 3(-1)^2 -2(-1)(0) + 11

OpenStudy (anonymous):

what from there though??

OpenStudy (mimi_x3):

Then use a calculator.

OpenStudy (anonymous):

i have to show work though...

myininaya (myininaya):

(-1)^2=?

OpenStudy (mimi_x3):

= 3 + 0 + 11 = 14 (-1)^2?? >_<

myininaya (myininaya):

(-1)^2=1 3(1)=3 -2(-1)(0)=0 so we have exactly what mimi has above

OpenStudy (mimi_x3):

lols, i thought i was wrong xD

myininaya (myininaya):

no i was trying to get c108 to answer some questions

OpenStudy (anonymous):

well the answer is supposed to be 14 but i dont get how.

OpenStudy (mimi_x3):

You don't get it >_< Just use a calculator then you will see. Infact, you can do it yourself, its easy just adding.

OpenStudy (anonymous):

yeah, i dont get it. because you aren't explaining it...

OpenStudy (mimi_x3):

Explanation isn't required. But I did explain to you, scroll up and see. Where are you stuck ? I dont get it.

OpenStudy (anonymous):

i get it thanks.